OpenStudy (anonymous):

Help.. Solve the system with the addition method: -a/4-b/4=1 -a/3+b/6=1

mathslover (mathslover):

\[-a/4-b/4=1\] \[-(a+b)/4=1\] \[-(a+b)=4\] \[a+b=-4\] \[b=-4-a\] put the value of b in 2nd equation : \[-a/3+b/6 = -a/3+(-4-a)\] \[-a/3-4-a\] \[(-a-3a-12)/3\] \[-4a-12)/3\] \[-4(a+3)/3\] \[-4(a+3)/3 = 1\] \[-4(a+3)=1\] \[a+3=-1/4\] \[a=-1/4-3\] \[a=-13/4\]

mathslover (mathslover):

b = -4-a b = -4+13/4 = -3/4

OpenStudy (anonymous):

for a it is -10/3 & for b it is -2/3

OpenStudy (anonymous):

Simplifying first equation : a + b = -4 Second Equation : -2a + b = 6 Subtracting first and second equation : 3a = -10 Therefore, a = -10/3.. Like this, b = -2/3 is the right answer...